OpenStudy (anonymous):

Suppose that x and y vary inversely and that x = 2 when y = 8. Write the function that models the inverse variation.

hartnn (hartnn):

Inverse variation equation is \[\large y=\frac{k}{x} \implies xy=k\] so, just plug in x=2, y=8 and find the value of k.

hartnn (hartnn):

can you find 'k' =.... ?

hartnn (hartnn):

@yourstruelya ?

hartnn (hartnn):

you just have to multiply x and y to get k. k= xy= 8*2 = ... ?

OpenStudy (anonymous):

16. lol sorry, I felt it open and walked in the other room.

hartnn (hartnn):

yes, so your equation is just xy=16 got it ?

hartnn (hartnn):

felt it open and walked in the other room again ?

hartnn (hartnn):

anyways, this is done, the function is xy=16. hope you got it.